OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. x2 - 20x + 100 = 36 A. {-6, 6} B. {-5/3, 5/3} C. {6, 10} D. {4, 16}

OpenStudy (hoblos):

x2 - 20x + 100 = 36 (x-10)^2 = 36 x-10 =±6 so x=16 or x=4

OpenStudy (hoblos):

D